OpenStudy (anonymous):

Choose the correct conic section to fit the equation. x2+ y2= 9

OpenStudy (anonymous):

anyone,,,its eaither ellipse circle parabola or hyperbola

OpenStudy (anonymous):

It's a equation of circle.

OpenStudy (anonymous):

look you can rewrite the equation like this (x-0)^2+(y-0)^2=3^2

OpenStudy (anonymous):

\[x ^{2}+y ^{2}=9, x ^{2}+y ^{2}=3^{2},It is a circle with centre \left( 0,0 \right) and radius =3\]